OpenStudy (gabby_taylor1224):

A cylinder has a radius of 4 feet and a height of 6 feet. Which of the following can be used to calculate the volume of a cone that fits exactly inside the cylinder? 1 over 3(3.14)(42)(6) (3.14)(42)(6) 1 over 3(3.14)(62)(4) (3.14)(62)(4

pooja195 (pooja195):

\[\huge~\rm~V=~\pi~r^2h/3\] This is the formula for the volume of a cone plug in your given values :)

pooja195 (pooja195):

\[\huge~\rm~V=~\pi~r^2\frac{ h }{ 3 }\]

pooja195 (pooja195):

Are you stuck?

OpenStudy (gabby_taylor1224):

I pulled the number in to the formulas and I got 100

OpenStudy (gabby_taylor1224):

plugged

pooja195 (pooja195):

We don't need to solve you need to plug in the given info into the formula

OpenStudy (gabby_taylor1224):

so would my answer be C?

pooja195 (pooja195):

Why do you feel that it is C?

OpenStudy (gabby_taylor1224):

I mean D because you are plugging in 3.14 and the radius

pooja195 (pooja195):

Im having trouble understanding your options \[\huge~\rm~ 1 /3(3.14)(4^2)(6)\] \[\huge~\rm~ (3.14)(4^2)(6)\] \[\huge~\rm~1/ 3(3.14)(6^2)(4)\] \[\huge~\rm~ (3.14)(6^2)(4)\]

pooja195 (pooja195):

Now try answering your question

pooja195 (pooja195):

C and D are incorrect

OpenStudy (gabby_taylor1224):

I would go with B then

pooja195 (pooja195):

Well done